### v2

# On the permutation automorphisms of binary cubic codes

Abstract.
We investigate the structural properties of binary cubic codes. We prove that up to dimension or codimension $4$, there is no binary linear code whose permutation automorphism group is generated by a fixed point free permutation of order $3$. We also prove that there is no $5$-dimensional binary code whose length is at least $30$ and whose permutation automorphism group is generated by a fixed point free permutation of order $3$. We also provide some computational results for the five dimensional binary cubic codes of length smaller than $30$.
Key words and phrases:
binary linear code, fixed point free permutation automorphism, fixed subcode associated to a permutation automorphism
2020 Mathematics Subject Classification:
94B05, 11T71, 20B05
1. Introduction
Let $C$ be a binary linear code and let $\mathrm{PAut}(C)$ denote its permutation automorphism group. Our motivation in this paper comes from the following question mentioned in  [3]:
Question 1.1.
Let $C$ be a (self-dual) binary linear code and suppose that $\mathrm{PAut}(C)$ contains a non-trivial subgroup $G$. What can we say about $C$ knowing $G$?
When ${\rm{PAut}}(C)$ is non-trivial, then for any subgroup $G\leq{\rm{PAut}}(C)$, $C$ becomes a module over the group algebra ${\mathbb{F}_{2}}G$. This leads us to study linear codes by means of different algebraic tools as done in [5] for instance. Additionally, the structure of $C$ can be determined by focusing on some specific automorphisms in ${\rm{PAut}}(C)$ (see [4, 8]). As a continuation of our study in [1], we study the properties of binary linear codes whose permutation automorphism group has a fixed point free permutation of order $3$. These codes are called binary cubic codes in the literature (see [2]). There is a famous problem in coding theory which asks the existence of a putative self-dual $[72,36,16]$-code and this problem comes down to the search of its permutation automorphism group for which the possibilities are $1,C_{2},C_{3},C_{2}\times C_{2},C_{5}$ (see [3] for a summary of this problem). We aim to get some insight and have a contribution to this problem in future by the methods that we invoke in this paper.
The main results of the paper are the following.
Theorem 1.2.
There is no binary cubic linear code of dimension $k$ where $k\in\{1,2,3,4\}$ having permutation automorphism group of order $3$.
Note that since a linear code has the same permutation automorphism group with its dual, by using this theorem we can deduce that there is no binary cubic linear code of length $n$, where $n$ is a multiple of $3$, and dimension $n-k$ where $k\in\{1,2,3,4\}$ having permutation automorphism group of order $3$.
Moreover, we have the following result for dimension $5$.
Theorem 1.3.
There is no binary $5$-dimensional cubic linear code of length $n\geq 30$ having permutation automorphism group of order $3$.
In [7], it is proved that all cubic codes can be obtained by a generalized cubic construction from a binary code and a quaternary code, both of the same length. Using the construction methods in [7] (see Theorem 5.1 and Section VI, case $m=3$ and Turyn’s Construction), we consider cubic codes of dimension $5$ and among them, we determine the smallest order for the permutation automorphism groups of cubic codes of length $n$ where $n\in\{6,9,12,15,18,21\}$ in SAGEMATH. For the length $n\in\{24,27\}$, one needs a very long time for the calculations in SAGEMATH, but for these lengths we have partial results which support the claim that a $5$-dimensional cubic code of length $n\in\{24,27\}$ has permutation automorphism group whose order is greater than $3$. (for detailed explanations see Remark 3.16).
The structure of the paper is as follows. In section 2, we introduce some notations and discuss lower dimensional cases. In section 3, we present the proofs of the main results.
2. Notation and results for lower dimensions
A binary linear code of length $n$ and dimension $k$ is a $k$-dimensional subspace of the vector space ${\mathbb{F}}_{2}^{n}$. The symmetric group $S_{n}$ acts on ${\mathbb{F}}_{2}^{n}$ as follows. For $\sigma\in S_{n}$ and for $v=v_{1}v_{2}\dots v_{n}$, we define $v^{\sigma}=v_{\sigma^{-1}(1)}v_{\sigma^{-1}(2)}\dots v_{\sigma^{-1}(n)}$. A permutation automorphism of $C$ is a permutation $\sigma\in S_{n}$ such that $C^{\sigma}=C$. The set of all automorphisms of $C$ forms a subgroup of $S_{n}$, which is called the permutation automorphism group of $C$ and it is denoted by ${\rm PAut}(C)$.
Let $C$ be a binary linear code. Assume that $\sigma$ is a fixed point free permutation of order $3$ in ${\rm PAut}(C)$. In this case, it is clear that $C$ has length a multiple of $3$. So for the rest of the paper, we assume that $C$ is a binary linear code of length $n=3m$ where $m$ is a positive integer. Since any conjugate of a permutation automorphism of $C$ is a permutation automorphism of a code which is permutation equivalent to $C$, without loss of generality, we can assume that
$\sigma=(1,2,3)(4,5,6)...(n-2,n-1,n).$
Let $\sigma_{i}$ be the $i^{th}$ cycle of $\sigma$ so that $\sigma=\prod_{i=1}^{m}\sigma_{i}$ and let $\Omega_{i}$ be the $i^{th}$ cycle set of $\sigma$ for $i=1,\ldots,m$. Given $v$ in $C$, let $v|_{\Omega_{i}}$ denote the $i$-th $3$-block of $v$ associated to the $i^{th}$ cycle set $\Omega_{i}$. The subcode $F_{\sigma}(C)=\{v\in C\ |\ v^{\sigma}=v\}$ is called the fixed subcode of $C$ associated to $\sigma$ and there is another subcode which is defined as follows
$E_{\sigma}(C)=\{v\in C\ |\ \sum_{i\in\Omega_{j}}v_{i}=0\text{ for any }j=1,\ldots m\}.$
It is clear that both $F_{\sigma}(C)$ and $E_{\sigma}(C)$ are $\sigma$-invariant.
The celebrated work of Cary Huffman gives a decomposition in terms of these subcodes associated to an automorphism.
Theorem 2.1.
[6, Theorem 1] With the notations above, we have that $C=F_{\sigma}(C)\oplus E_{\sigma}(C)$.
Moreover, Huffman proves the following result which plays an important role in our discussion.
Lemma 2.2.
[6, Corollary 1]  Let $C$ be a linear code whose base field is $\mathbb{F}_{q}$ and let $\sigma$ be a permutation in ${\rm PAut}(C)$. If the order of $\sigma$ is a power of a prime number $s$, then the multiplicative order of $q$ modulo $s$, divides $\mathrm{dim}(E_{\sigma}(C))$.
Let $C$ be a linear code whose permutation automorphism group contains a fixed point free permutation $\sigma$ of order $3$. Such a code is called cubic (see [2, 7]). Moreover, Lemma  2.2 implies that $2$ divides ${\rm dim}(E_{\sigma}(C))$. We start our discussion introducing some notations which will be used in our proofs.
Given a non-zero vector $v$ in $E_{\sigma}(C)$, assume that $v|_{\Omega_{i}}$ is a non-zero $3$-block of $v$, then it follows from the definition of $E_{\sigma}(C)$ that there are exactly two $1$’s in this block. Let $a_{i,v,1}$, $a_{i,v,2}$, and $a_{i,v,3}$ denote the positions of the first $1$, the second $1$, and zero in the $3$-block $v|_{\Omega_{i}}$, respectively. Moreover, let ${\rm J}_{v}$ be the set of indices $i$ for which $v|_{\Omega_{i}}$ is a non-zero $3$-block. Assume $0\neq w\in E_{\sigma}(C)$ is a vector different from $v$ whose weight is equal to that of $v$. If $wt(v)=wt(w)=2m$, then this implies that neither $v$ nor $w$ has a zero $3$-block. If $wt(v)=wt(w)=2t<2m$, observe that $|{\rm J}_{v}\setminus{\rm J}_{w}|=|{\rm J}_{w}\setminus{\rm J}_{v}|$. Let $d:=|{\rm J}_{v}\setminus{\rm J}_{w}|=|{\rm J}_{w}\setminus{\rm J}_{v}|$. If ${\rm J}_{v}\setminus{\rm J}_{w}\neq\emptyset$, then $d\geq 1$. Moreover if $i\in{\rm J}_{v}\setminus{\rm J}_{w}$, then $v|_{\Omega_{i}}$ is a non-zero $3$-block while $w|_{\Omega_{i}}$ is a zero $3$-block. Let
${\rm J}_{v}\setminus{\rm J}_{w}:=\{r_{w,1},...,r_{w,d}\}$
and
${\rm J}_{w}\setminus{\rm J}_{v}:=\{r_{v,1},...,r_{v,d}\}$
where for each $j\in\{1,...,d\}$, the symbols $r_{v,j}$ denote the $j$-th zero $3$-block in $v$ whose associated $3$-block in $w$ is non-zero.
Example 2.3.
Let $\sigma=(1,2,3)(4,5,6)(7,8,9)(10,11,12)(13,14,15)(16,17,18)$ and $C$ a binary linear code spanned by the following vectors:
[…]
It is clear that $wt(v)=wt(w)=6<2m=12$. Moreover, we have that $\sigma\in{\rm PAut}(C)$ and $C=E_{\sigma}(C)$. On the other hand $J_{v}\backslash J_{w}=\{1,4\}$, $J_{w}\backslash J_{v}=\{2,6\}$, $r_{w,1}=1$, $r_{w,2}=4$, $a_{r_{w,1},v,1}=a_{1,v,1}=1$, and $a_{r_{v,1},w,1}=a_{2,w,1}=4$.
We start our investigation with the simplest case for completeness. If $C$ is a $1$-dimensional binary linear code of length $n$, then $\mathrm{PAut}(C)$ is not generated by a fixed point free permutation of order $3$. Because if $v$ is the non-zero vector of $C$, then $\mathrm{PAut}(C)$ is equal to stabilizer of $v$ which is isomorphic to $S_{t}\times S_{n-t}$ where $t$ is equal to the weight of $v$.
Lemma 2.4.
Let $C$ be a binary linear code and $\sigma$ a fixed point free automorphism of $C$ of order $3$. Then for any $v\in E_{\sigma}(C)$, we have that $v^{\sigma^{2}}+v^{\sigma}+v=0$.
Proof.
We claim that $(v^{\sigma^{2}}+v^{\sigma}+v)|_{\Omega_{j}}=000$ for each $j=1,2,\ldots,m$. Since $C$ is a binary code, the definition of $E_{\sigma}(C)$ implies that $v|_{\Omega_{j}}$ is either a zero $3$-block or it has exactly two $1$’s. If it is a zero $3$-block, our claim holds trivially. If $v|_{\Omega_{j}}$ has exactly two $1^{\prime}s$, then in every position of the block $\Omega_{j}$, exactly two of the vectors $v^{\sigma^{2}}$, $v^{\sigma}$ and $v$ have value equal to $1$. So the claim holds and the proof is finished since $\sigma$ is fixed point free. ∎
Theorem 2.5.
Let $C$ be a $k$-dimensional binary linear code where $k\geq 2$ and let $\sigma$ be a fixed point free permutation of order $3$ in ${\rm PAut}(C)$. If ${\rm dim}(E_{\sigma}(C))=2$, then there exists an involution $\theta$ in ${\rm PAut}(E_{\sigma}(C))$ which fixes every element of $F_{\sigma}(C)$. Indeed, $\theta\in{\rm PAut}(C)$.
Proof.
Let $v$ be a non-zero vector in $E_{\sigma}(C)$. Since $v^{\sigma}\neq v$ and $E_{\sigma}(C)$ is $\sigma$-invariant, letting $w:=v^{\sigma}$, we deduce that $\{v,w\}$ is a basis for $E_{\sigma}(C)$.
First assume that $v$ has no zero $3$-block. Since $\sigma=(1,2,3)\ldots(n-2,n-1,n)$, and $v\in E_{\sigma}(C)$, in each $3$-block of $v$ we have a unique zero.
Let $\alpha_{i}=(a_{i,v,3},a_{i,w,3})$, where $a_{i,v,3}$ denotes the position of the unique zero in the $3$-block $v|_{\Omega_{i}}$, for each $i\in\{1,...,m\}$. Define $\theta:=\Pi_{i=1}^{m}\alpha_{i}$. The fact that $(v|_{\Omega_{i}})^{\alpha_{i}}=w|_{\Omega_{i}}$ for each $i\in\{1,...,m\}$ implies that $v^{\theta}=w$. In particular, ${\rm PAut}(E_{\sigma}(C))$ has an involution. If $C=E_{\sigma}(C)$, we are done with the proof, so assume that ${\rm dim}(F_{\sigma}(C))\geq 1$. Let $z$ be an arbitrary non-zero vector in $F_{\sigma}(C)$. The definition of $F_{\sigma}(C)$ verifies that all $3$-blocks of $z$ are either $000$ or $111$. Since, $\theta$ stabilizes these $3$-blocks, we deduce that $z^{\theta}=z$ for any $z\in F_{\sigma}(C)$. Hence $\theta\in{\rm PAut}(C)$.
It should be mentioned that $v$ has a zero $3$-block if and only if $v^{\sigma}$ and $v^{\sigma^{2}}$ have a zero $3$-block at the same position. So it is enough to concentrate on non-zero $3$-blocks of $v$ and apply the previous argument to non-zero blocks of $v$ in order to obtain an involution in $\mathrm{PAut}(C)$.
Corollary 2.6.
Let $k\in\{2,3\}$. There exists no $k$-dimensional binary linear code $C$ which satisfies ${\mathrm{PAut}}(C)=\langle\sigma\rangle$ where $\sigma$ is a fixed point free permutation of order $3$.
Proof.
Notice that if $C$ is a binary linear code such that $C=F_{\sigma}(C)$, then the case $\mathrm{PAut}(C)=\langle\sigma\rangle$ does not occur. The rest is just a consequence of Lemma  2.2 and Theorem  2.5.
Corollary 2.7.
Let $C$ be a $k$-dimensional binary linear code where ${\rm PAut}(C)=\langle\sigma\rangle$ and $\sigma$ is a fixed point free permutation of order $3$. Then $k\geq 4$ and ${\rm dim}(E_{\sigma}(C))\geq 4$.
Proof.
Due to Corollary  2.6 we come to the conclusion that $k\geq 4$. Moreover, Lemma  2.2 verifies that $2$ divides $\mathrm{dim}(E_{\sigma}(C))$. Now Theorem  2.5 guarantees that ${\rm dim}(E_{\sigma}(C))\neq 2$, hence ${\rm dim}(E_{\sigma}(C))\geq 4$.
3. Proof of the main results
We start with recording the following elementary fact.
Lemma 3.1.
Let $C$ be a binary linear code and $\sigma$ a fixed point free permutation of order $3$ in ${\rm PAut}(C)$. Let $D$ be a $2r$-dimensional $\sigma$-invariant subspace of $E_{\sigma}(C)$. Then a basis of $D$ can be chosen as $\beta=\{v_{(1)},...,v_{(r)},v_{(1)}^{\sigma},...,v_{(r)}^{\sigma}\}$, where $v_{(i+1)}\notin\langle v_{(1)},...,v_{(i)},v_{(1)}^{\sigma},...,v_{(i)}^{\sigma}\rangle$, for each $i\in\{1,...,r-1\}$.
Proof.
We first claim that $V:=\langle v_{(1)},...,v_{(r)},v_{(1)}^{\sigma},...,v_{(r)}^{\sigma}\rangle$ is a $\sigma$-invariant vector space. Let $y\in V$, then
$y=(\sum_{i=1}^{r}a_{i}v_{(i)})+(\sum_{i=1}^{r}b_{i}v_{(i)}^{\sigma})$
for some $a_{i},b_{i}\in{\mathbb{F}}_{2}$. It follows that
$y^{\sigma}=(\sum_{i=1}^{r}a_{i}v_{(i)}^{\sigma})+(\sum_{i=1}^{r}b_{i}v_{(i)}^{\sigma^{2}}).$
But by Lemma 2.4, $v_{(i)}^{\sigma^{2}}=v_{(i)}+v_{(i)}^{\sigma}$, hence we get that
$y^{\sigma}=(\sum_{i=1}^{r}b_{i}v_{(i)})+(\sum_{i=1}^{r}(a_{i}+b_{i})v_{(i)}^{\sigma}),$
so the claim follows.
We next claim that $\{v_{(1)},...,v_{(r)},v_{(1)}^{\sigma},...,v_{(r)}^{\sigma}\}$ is a linearly independent set. This is clear for $r=1$ since $v\neq v^{\sigma}$ for any non-zero $v\in E_{\sigma}(C)$. For $r>1$, let
$(\sum_{i=1}^{r}a_{i}v_{(i)})+(\sum_{i=1}^{r}b_{i}v_{(i)}^{\sigma})=0$
for $a_{i},b_{i}\in{\mathbb{F}}_{2}$. Assume at least one of $a_{r}$ or $b_{r}$ is non-zero. Then we have that
$a_{r}v_{(r)}+b_{r}v_{(r)}^{\sigma}=(\sum_{i=1}^{r-1}a_{i}v_{(i)})+(\sum_{i=1}^{r-1}b_{i}v_{(i)}^{\sigma}).\quad(**)$
Let’s analyse possible cases. If $a_{r}=1$ and $b_{r}=0$, then ()(**)( ∗ ∗ ) implies that
$v_{(r)}\in\langle v_{(1)},...,v_{(r-1)},v_{(1)}^{\sigma},...,v_{(r-1)}^{\sigma}\rangle$
a contradiction. If $a_{r}=0$ and $b_{r}=1$, then ()(**)( ∗ ∗ ) implies that
$v_{(r)}^{\sigma}\in\langle v_{(1)},...,v_{(r-1)},v_{(1)}^{\sigma},...,v_{(r-1)}^{\sigma}\rangle$
but by the previous paragraph since $\langle v_{(1)},...,v_{(r-1)},v_{(1)}^{\sigma},...,v_{(r-1)}^{\sigma}\rangle$ is $\sigma$-invariant, this is impossible. If $a_{r}=1$ and $b_{r}=1$, then ()(**)( ∗ ∗ ) implies that
$v_{(r)}^{\sigma^{2}}=v_{(r)}+v_{(r)}^{\sigma}\in\langle v_{(1)},...,v_{(r-1)},v_{(1)}^{\sigma},...,v_{(r-1)}^{\sigma}\rangle$
but again $\sigma$-invariance gives a contradiction. Hence the claim is established.
Now choose a non-zero vector $v\in D$ and set $v_{(1)}:=v$. Then choose $v_{(2)}\in D$ such that $v_{(2)}\notin\langle v_{(1)},v_{(1)}^{\sigma}\rangle$ and continue this process until $D=\langle v_{(1)},...,v_{(r)},v_{(1)}^{\sigma},...,v_{(r)}^{\sigma}\rangle$. This finishes the proof.
For a permutation automorphism $\sigma$ of $\mathrm{PAut}(C)$, it is easy to see that all codewords lying in the same orbit of $\langle\sigma\rangle$ have equal weights.
Definition 3.2.
Let $C$ be a binary linear code and $\sigma\in{\rm PAut}(C)$. $C$ is called a distinct $\sigma$-weight code if any two $\langle\sigma\rangle$-orbits of $C$ have different weights. Otherwise, $C$ is called a non-distinct $\sigma$-weight code.
Example 3.3.
Let $C$ be a binary linear code spanned by the following vectors:
$v_{1}=111000,\ v_{2}=101101,\ v_{3}=110110.$
Using ${\rm MAGMA}$, we find out that $|{\rm PAut}(C)|=2^{3}\cdot 3$. Indeed, $\sigma=(1,2,3)(4,5,6)$ is in $\mathrm{PAut}(C)$. Moreover, $C$ is a non-distinct $\sigma$-weight code as $v_{1}$ and ${v_{1}+v_{2}}$ are in two different $\langle\sigma\rangle$-orbits but their weights are equal.
Proposition 3.4.
Let $C$ be a binary linear code such that ${\rm PAut}(C)$ contains $\sigma$ which is a fixed point free permutation of order $3$. If $E_{\sigma}(C)$ is a $4$-dimensional distinct $\sigma$-weight code, then the length of $C$ is at least $30$.
Proof.
Since $4={\rm dim}(E_{\sigma}(C))$, Lemma  3.1 guarantees the existence of two linearly independent vectors $v$ and $w$ in $E_{\sigma}(C)$ such that $\beta=\{v,w,v^{\sigma},w^{\sigma}\}$ is a basis for $E_{\sigma}(C)$. Let $l$ denote the number of $1$’s which overlap in $v$ and $w$. Then it is not hard to observe that
$wt(v+w)=wt(w)+wt(w)-2l.$
For $i\in\{0,1,2\}$, let
$l_{i}:=|\{j\in{\rm J}_{v^{\sigma^{i}}}\cap{\rm J}_{w}:v^{\sigma^{i}}|_{\Omega_{j}}=w|_{\Omega_{j}}\}|$
denote the number of non-zero $3$-blocks such that $v^{\sigma^{i}}$ and $w$ coincide. Let $s:=|{\rm J}_{w}\setminus{\rm J}_{v}|$ denote the number of non-zero $3$-blocks of $w$ such that the associated $3$-block of $v$ is zero. In a similar way, let $t:=|{\rm J}_{v}\setminus{\rm J}_{w}|$. Let $r:=|\{j\notin{\rm J}_{v}\cup{\rm J}_{w}\}|$ be the number of common zero $3$-blocks of $v$ and $w$. One can observe that
$n=3(l_{0}+l_{1}+l_{2}+s+r+t)$
$wt(v)=2(l_{0}+l_{1}+l_{2})+2t$
$wt(w)=2(l_{0}+l_{1}+l_{2})+2s$
$wt(v+w)=2(l_{1}+l_{2}+t+s)$
$wt(v^{\sigma}+w)=2(l_{0}+l_{2}+t+s)$
$wt(v^{\sigma^{2}}+w)=2(l_{0}+l_{1}+t+s).$
Moreover, as these vectors belong to distinct orbits and due to the hypothesis that $E_{\sigma}(C)$ is a distinct $\sigma$-weight code, we deduce that $\{l_{0},l_{1},l_{2},s,t\}$ is a set which contains five distinct non-negative integers. Hence the minimal choice would be $\{0,1,2,3,4\}$ which forces $n$ be at least $30$. ∎
The following example shows that the lower bound given in Proposition  3.4 for the length of such codes is attained.
Example 3.5.
Let $C$ be a binary linear code spanned by
[…]
Using ${\rm MAGMA}$, we calculate that $|{\rm PAut}(C)|=2^{15}\cdot 3^{7}$. In particular,
$\sigma=\prod_{i=0}^{9}(3i+1,3i+2,3i+3)$
is in ${\rm PAut}(C)$ and $C=E_{\sigma}(C)$ is a $4$-dimensional binary linear code with length $30$. Moreover, we observe that $C$ is a distinct $\sigma$-weight code.
Using the notations mentioned in Proposition  3.4, we state the following result.
Corollary 3.6.
Let $C$ be a binary linear code where ${\rm PAut}(C)$ contains $\sigma$ which is a fixed point free permutation of order $3$. If $E_{\sigma}(C)$ is a $4$-dimensional distinct $\sigma$-weight code, then $l_{i}\geq 2$ for some $i\in\{0,1,2\}$.
Proof.
This follows from the proof of Proposition  3.4
Theorem 3.7.
Let $\sigma$ be a fixed point free permutation of order $3$ in $S_{n}$. There is no binary $4$-dimensional non-distinct $\sigma$-weight code $C$ where ${\rm PAut}(C)$ is generated by $\sigma$.
Proof.
Suppose that $C$ is a binary $4$-dimensional non-distinct $\sigma$-weight code where $\mathrm{PAut}(C)$ is generated by $\sigma$. First of all, Corollary  2.7 implies that $C=E_{\sigma}(C)$. Since $4={\rm dim}(E_{\sigma}(C))={\rm dim}(C)$, Lemma  3.1 guarantees the existence of two linearly independent vectors $v$ and $w$ in $E_{\sigma}(C)$ such that $\beta=\{v,w,v^{\sigma},w^{\sigma}\}$ is a basis for $E_{\sigma}(C)$. Since $C$ is a non-distinct $\sigma$-weight code, without loss of generality, we may assume that $wt(v)=wt(w)$. We have the following cases.
Case 1. Assume that $wt(v)=wt(w)=2m$, where $n=3m$ is the length of $C$. This implies that neither $v$ nor $w$ has a zero $3$-block. For $i\in\{1,...,m\}$, if $v|_{\Omega_{i}}\neq w|_{\Omega_{i}}$, let $\alpha_{i}$ be the transposition with $(v|_{\Omega_{i}})^{\alpha_{i}}=w|_{\Omega_{i}}$ which is the transposition that interchanges the positions of the unique zero in $v|_{\Omega_{i}}$ and $w|_{\Omega_{i}}$. If $v|_{\Omega_{i}}=w|_{\Omega_{i}}$, let $\alpha_{i}$ be an arbitrary transposition associated to the elements in that block. Let $\alpha:=\prod_{i=1}^{m}\alpha_{i}$. Then it is easy to see that $v^{\alpha}=w$. Moreover, since $\sigma=\prod_{i=1}^{m}\sigma_{i}$ and each $i$ satisfies $\alpha_{i}^{-1}\sigma_{i}\alpha_{i}=\sigma_{i}^{-1}$ we have that $\alpha^{-1}\sigma\alpha=\sigma^{-1}$, so we have that
$(v^{\sigma})^{\alpha}=w^{\sigma^{2}}\text{ and }(v^{\sigma^{2}})^{\alpha}=w^{\sigma}.$
Therefore, $\alpha\in{\rm PAut}(C)$.
Case 2. Assume that $wt(v)=wt(w)=2t<2m$. From Section 2, we know that $|{\rm J}_{v}\setminus{\rm J}_{w}|=|{\rm J}_{w}\setminus{\rm J}_{v}|$. We will construct a permutation automorphism of $C$ different from $\sigma$ via the following steps.
  • (1)
    If $i\in{\rm J}_{v}\cap{\rm J}_{w}$, then $v|_{\Omega_{i}}$ and $w|_{\Omega_{i}}$ are both non-zero $3$-blocks. In this case, we can construct an involution $\alpha_{i}$ as we did in Case 1.
  • (2)
    If $i\notin{\rm J}_{v}\cup{\rm J}_{w}$, then $v|_{\Omega_{i}}$ and $w|_{\Omega_{i}}$ are both zero $3$-blocks. In this case, let $\alpha_{i}$ to be an arbitrary transposition associated to the elements in block $i$.
  • (3)
    If $i\in{\rm J}_{v}\setminus{\rm J}_{w}\neq\emptyset$. Recall that $|{\rm J}_{v}\setminus{\rm J}_{w}|=|{\rm J}_{w}\setminus{\rm J}_{v}|$. Letting $d=|{\rm J}_{v}\setminus{\rm J}_{w}|$, then $d\geq 1$.
    Let
    ${\rm J}_{v}\setminus{\rm J}_{w}:=\{r_{w,1},...,r_{w,d}\}\text{ and }{\rm J}_{w}\setminus{\rm J}_{v}:=\{r_{v,1},...,r_{v,d}\}.$
    We construct an involution which in particular, interchanges the $v|_{\Omega_{r_{w,j}}}$ and $w|_{\Omega_{r_{v,j}}}$ for each $j\in\{1,\ldots,d\}$. Namely, let $\hat{\alpha}_{j}$ send the position of zero in the block $r_{w,j}$ of $v$ to the position of zero in the block $r_{v,j}$ of $w$, and adjust the other positions of these blocks so that $\hat{\alpha}_{j}$ satisfies
    $\hat{\alpha}_{j}^{-1}\sigma_{r_{w,j}}\sigma_{r_{v,j}}\hat{\alpha}_{j}=\sigma_{r_{v,j}}^{-1}\sigma_{r_{w,j}}^{-1}.$
Let us define
$\alpha=\prod_{i\in{\rm J}_{v}\cap{\rm J}_{w}}\alpha_{i}\prod_{i\notin{\rm J}_{v}\cup{\rm J}_{w}}\alpha_{i}\prod_{j\in\{1,...,d\}}\hat{\alpha}_{j}.$
It is easy to see that $v^{\alpha}=w$. Moreover, $\alpha$ satisfies $\alpha^{-1}\sigma\alpha=\sigma^{-1}$. As a result, we have the following equalities:
$(v^{\sigma})^{\alpha}=(v^{\alpha})^{\sigma^{2}}=w^{\sigma^{2}}$
$(w^{\sigma})^{\alpha}=(w^{\alpha})^{\sigma^{2}}=v^{\sigma^{2}}$
Hence we conclude that $\alpha\in{\rm PAut}(E_{\sigma}(C))={\rm PAut}(C)$, a contradiction. ∎
Example 3.8.
Let $C$ be the code defined in Example 2.3. Recall that,
$\sigma=(1,2,3)(4,5,6)(7,8,9)(10,11,12)(13,14,15)(16,17,18)$
and $C$ is a binary linear code spanned by
[…]
We have that $J_{v}\cap J_{w}=\{3\},\ (J_{v}\cup J_{w})^{\prime}=\{5\},\ J_{v}\backslash J_{w}=\{1,4\},\ J_{w}\backslash J_{v}=\{2,6\}$. Moreoever,
$r_{w,1}=1,\ r_{w,2}=4\text{ and }r_{v,1}=2,\ r_{v,2}=6,$
which gives us that $\alpha_{3}=(8,9),\alpha_{5}=(13,14)$, $\hat{\alpha}_{1}=(1,4)(2,6)(3,5)$, $\hat{\alpha}_{2}=(11,16)(12,18)(10,17)$ and it follows that
$\alpha=(1,4)(2,6)(3,5)(8,9)(10,17)(11,16)(12,18)(13,14).$
We observe that $v^{\alpha}=w$ and
$\alpha^{-1}\sigma\alpha=(4,6,5)(1,3,2)(7,9,8)(17,16,18)(14,13,15)(11,10,12)=\sigma^{-1}.$
Corollary 3.9.
Let $\sigma$ be a fixed point free permutation of order $3$. If $C$ is a $4$-dimensional non-distinct $\sigma$-weight code, then $S_{3}\leq\mathrm{PAut}(C)$.
Proof.
Follows from the proof of Theorem 3.7. ∎
Proposition 3.10.
Let $C$ be a binary linear code and $\sigma$ a fixed point free permutation of order $3$ in ${\rm PAut}(C)$. If $C$ is a $4$-dimensional distinct $\sigma$-weight code and $C=E_{\sigma}(C)$, then $\mathrm{PAut}(C)$ contains an involution.
Proof.
As a result of Lemma  3.1 and due to the hypothesis, there exists a basis $\beta=\{v,w,v^{\sigma},w^{\sigma}\}$ for $E_{\sigma}(C)$ where $wt(v)\neq wt(w)$. Without loss of generality, we may assume that $v$ has some zero $3$-blocks.
Case 1. If $v$ has at least two zero $3$-blocks, then for some distinct $i$ and $j$, we have $v|_{\Omega_{i}}=v|_{\Omega_{j}}=000$. Now we have the following possibilities.
  • (1)
    $w|_{\Omega_{i}},w|_{\Omega_{j}}\in\{110,011,101\}$ or $w|_{\Omega_{i}}=w|_{\Omega_{j}}=000$. In any case, there exist some positions $e$ and $f$ in the $i^{th}$ and the $j^{th}$ $3$-blocks, respectively, for which $w_{e}=w_{f}=0$. Set $\alpha=(e,f)$. Now, we can see that
    $v^{\alpha}=v,\quad(v^{\sigma})^{\alpha}=v^{\sigma},\quad(v^{\sigma^{2}})^{\alpha}=v^{\sigma^{2}}$
    Moreover, the equalities $(w^{\sigma})_{e}=(w^{\sigma})_{f}=1$ and $(w^{\sigma^{2}})_{e}=(w^{\sigma^{2}})_{f}=1$ verify that
    $w^{\alpha}=w,\quad(w^{\sigma})^{\alpha}=w^{\sigma},\quad(w^{\sigma^{2}})^{\alpha}=w^{\sigma^{2}}$
    Hence $\alpha\in\mathrm{PAut}(C)$.
  • (2)
    $w|_{\Omega_{i}}\in\{110,011,101\}$ and $w|_{\Omega_{j}}=000$. Focus on the block $\Omega_{j}$. Since, for any $e$ and $f$ in $\Omega_{j}$ satisfy $w_{e}=w_{f}=0$ and $v_{e}=v_{f}=0$, we can observe that the transposition $\alpha:=(e,f)$ fixes every element of $C$. Hence $\alpha\in\mathrm{PAut}(C)$.
Case 2. If $v$ has a unique zero $3$-block, then Corollary  3.6 verifies that at least one of the elements of $\{l_{0},l_{1},l_{2}\}$ must be greater than or equal to $2$. Without loss of generality, assume that $l_{0}\geq 2$. This implies that ${\rm J}_{v}\cap{\rm J}_{w}$ includes at least two distinct elements, say $i$ and $j$, for which $v|_{\Omega_{i}}=w|_{\Omega_{i}}$ and $v|_{\Omega_{j}}=w|_{\Omega_{j}}$. As a result, $(v+w)|_{\Omega_{i}}=(v+w)|_{\Omega_{j}}=000$. Now $v+w$ and $w$ satisfy the conditions in (1) of Case 1. ∎
Corollary 3.11.
There exists no $4$-dimensional binary linear code $C$ such that ${\rm PAut}(C)=\langle\sigma\rangle$ where $\sigma$ is a fixed point free permutation of order three.
Proof.
This is a direct consequence of Corollary  2.7, Theorem  3.7, and Proposition  3.10.
Corollary 3.12.
Let $C$ be a binary linear code and $\mathrm{PAut}(C)=\langle\sigma\rangle$ where $\sigma$ is a fixed point free permuation of order $3$. If $C=E_{\sigma}(C)$, then $\mathrm{dim}(C)\geq 6$.
Proposition 3.13.
Let $\sigma$ be a fixed point free permutation of order $3$. There is no binary $5$-dimensional linear code $C$ such that $E_{\sigma}(C)$ is distinct $\sigma$-weight code and ${\rm PAut}(C)$ is generated by $\sigma$.
Proof.
Suppose that $C$ is a $5$-dimensional linear code such that ${\rm PAut}(C)=\langle\sigma\rangle$ and $E_{\sigma}(C)$ is a distinct $\sigma$-weight code. Then by Proposition 3.4 we have that the length of $C$ is at least $30$ and by Corollary 2.7, we have that ${\rm dim}(E_{\sigma}(C))=4$. On the other hand, Lemma 3.1 implies that we have a basis $\{v,w,v^{\sigma},w^{\sigma}\}$ for $E_{\sigma}(C)$. Let $z$ be the basis element in $F_{\sigma}(C)$. We use the notations in the proof of Proposition 3.4. Recall that $s:=|{\rm J}_{w}\setminus{\rm J}_{v}|$, $t:=|{\rm J}_{v}\setminus{\rm J}_{w}|$ and $r:=|\{j\notin{\rm J}_{v}\cup{\rm J}_{w}\}|$ where ${\rm J}_{v}$ denotes the set of indices $i$ for which $v|_{\Omega_{i}}$ is a non-zero $3$-block. For $i\in\{0,1,2\}$, let $l_{i}:=|\{j\in{\rm J}_{v^{\sigma^{i}}}\cap{\rm J}_{w}:v^{\sigma^{i}}|_{\Omega_{j}}=w|_{\Omega_{j}}\}|$ denote the number of non-zero $3$-blocks such that $v^{\sigma^{i}}$ and $w$ coincide. We have $n=3(l_{0}+l_{1}+l_{2}+s+r+t)$. As $E_{\sigma}(C)$ is a distinct $\sigma$-weight code, by Corollary 3.6, we have $l_{i}\geq 2$ for some $i\in\{0,1,2\}$. We have two cases:
Case 1. If $l_{i}\geq 3$ for some $i\in\{0,1,2\}$, without loss of generality, we can assume that $l_{0}\geq 3$. Since $l_{0}\geq 3$, there exists $i,j,k$ such that $v|_{\Omega_{i}}=w|_{\Omega_{i}},v|_{\Omega_{j}}=w|_{\Omega_{j}},v|_{\Omega_{k}}=w|_{\Omega_{k}}$. Consider the corresponding blocks $z|_{\Omega_{i}},z|_{\Omega_{j}},z|_{\Omega_{k}}$. As $z\in F_{\sigma}(C)$, at least two of the blocks $z|_{\Omega_{i}},z|_{\Omega_{j}},z|_{\Omega_{k}}$ are identical. Suppose $z|_{\Omega_{i}}=z|_{\Omega_{j}}$. Consider the generator matrix $G$ of $C$.
$G=\begin{bmatrix}z\\ v\\ w\\ v^{\sigma}\\ w^{\sigma}\end{bmatrix}=\begin{bmatrix}...&...&z|_{\Omega_{i}}&\dots&z|_{\Omega_{j}}&...\\ ...&....&v|_{\Omega_{i}}&\dots&v|_{\Omega_{j}}&...\\ ...&...&w|_{\Omega_{i}}&\ddots&w|_{\Omega_{j}}&...\\ ...&....&v^{\sigma}|_{\Omega_{i}}&\dots&v^{\sigma}|_{\Omega_{j}}&...\\ ...&...&w^{\sigma}|_{\Omega_{i}}&\ddots&w^{\sigma}|_{\Omega_{j}}&...\par\par\end{bmatrix}$
As $z|_{\Omega_{i}}=z|_{\Omega_{j}}$ and $v|_{\Omega_{i}},v|_{\Omega_{j}},w|_{\Omega_{i}},w|_{\Omega_{j}},v^{\sigma}|_{\Omega_{i}},w^{\sigma}|_{\Omega_{i}}\in\{110,011,101\}$ the following submatrix of $G$
$\begin{bmatrix}z|_{\Omega_{i}}&z|_{\Omega_{j}}&\\ v|_{\Omega_{i}}&v|_{\Omega_{j}}&\\ w|_{\Omega_{i}}&w|_{\Omega_{j}}&\\ v^{\sigma}|_{\Omega_{i}}&v^{\sigma}|_{\Omega_{j}}&\\ w^{\sigma}|_{\Omega_{i}}&w^{\sigma}|_{\Omega_{j}}&\end{bmatrix}$
has identical columns in pairs. Indeed, let $e$ be the position of zero in $v|_{\Omega_{i}}$ that is $v_{e}=0$ and let $f$ be the position of zero in $v|_{\Omega_{j}}$, that is $v_{f}=0$. Then $w_{e}=0=w_{f}$ and $v^{\sigma}_{e}=v^{\sigma}_{f}=w^{\sigma}_{e}=w^{\sigma}_{f}=1$. Let $\alpha=(e,f)$. In this case we have $z^{\alpha}=z,v^{\alpha}=v,w^{\alpha}=w,(v^{\sigma})^{\alpha}=v^{\sigma},(w^{\sigma})^{\alpha}=w^{\sigma}$ Hence $\alpha\in\mathrm{PAut}(C)$.
Case 2. Assume for all $i\in\{0,1,2\}$ $l_{i}\leq 2$, that is $\{l_{0},l_{1},l_{2}\}=\{0,1,2\}$. Since $E_{\sigma}(C)$ is a distinct $\sigma$-weight code, from the proof of Proposition 3.4, we get that $s$ or $t$ is greater than or equal to $3$. That means $v$ or $w$ has at least three zero blocks. Without loss of generality assume $w$ has three zero blocks, $w|_{\Omega_{i}},w|_{\Omega_{j}},w|_{\Omega_{k}}$. Then a submatrix of the generator matrix $G$ is of the form
$\begin{bmatrix}z|_{\Omega_{i}}&z|_{\Omega_{j}}&z|_{\Omega_{k}}\\ v|_{\Omega_{i}}&v|_{\Omega_{j}}&v|_{\Omega_{k}}\\ 000&000&000\\ v^{\sigma}|_{\Omega_{i}}&v^{\sigma}|_{\Omega_{j}}&v^{\sigma}|_{\Omega_{k}}\\ 000&000&000\end{bmatrix}.$
By definition of $s$ and $t$, all blocks $v|_{\Omega_{i}},v|_{\Omega_{j}},v|_{\Omega_{k}}$ are non-zero. Moreover, since $z\in F_{\sigma}(C)$, at least two of the blocks $z|_{\Omega_{i}},z|_{\Omega_{j}},z|_{\Omega_{k}}$ are identical, say $z|_{\Omega_{i}}=z|_{\Omega_{j}}$. In this case, we have a submatrix of $G$
$\begin{bmatrix}aaa&aaa&\\ v|_{\Omega_{i}}&v|_{\Omega_{j}}&\\ 000&000&\\ v^{\sigma}|_{\Omega_{i}}&v^{\sigma}|_{\Omega_{j}}&\\ 000&000&\end{bmatrix}$
where $a\in\{0,1\}$. As $v|_{\Omega_{i}},v|_{\Omega_{j}},v^{\sigma}|_{\Omega_{i}},v^{\sigma}|_{\Omega_{j}}\in\{110,011,101\}$, there are some columns of this submatrix which are identical in pairs. Indeed, let $e$ be the position of zero in the block $v|_{\Omega_{i}}$, that is $v_{e}=0$ and let $f$ be the position of zero in $v|_{\Omega_{j}}$, that is $v_{f}=0$. Then $v^{\sigma}_{e}=1=v^{\sigma}_{f}$. Let $\alpha=(e,f)$. It is clear that $\alpha$ fixes all basis elements and hence $\alpha\in\mathrm{PAut}(C)$.
Since in all possible cases, we deduce that there is an involution in $\mathrm{PAut}(C)$, the proof is finished. ∎
Proposition 3.14.
Let $\sigma$ be a fixed point free permutation of order $3$ in $S_{n}$. There is no binary $5$-dimensional linear code $C$ of length $n\geq 30$ such that $E_{\sigma}(C)$ is a non-distinct $\sigma$-weight code and ${\rm PAut}(C)$ is generated by $\sigma$.
Proof.
Let $\sigma$ be a fixed point free permutation of order $3$ in $S_{n}$ where $n\geq 30$. Suppose that $C$ is a $5$-dimensional linear code such that ${\rm PAut}(C)=\langle\sigma\rangle$ and $E_{\sigma}(C)$ is non-distinct $\sigma$-weight code. By Corollary 2.7 and by Lemma 3.1, we have a basis $\{v,w,v^{\sigma},w^{\sigma}\}$ for $E_{\sigma}(C)$. Since $E_{\sigma}(C)$ is a non-distinct $\sigma$-weight code, without loss of generality we can assume that $wt(v)=wt(w)$. Let $z$ be the basis element for $F_{\sigma}(C)$. Let $l_{0},l_{1},l_{2},r,s,t$ be the set of integers as defined in the proof of Proposition 3.4. Recall that $n=3(l_{0}+l_{1}+l_{2}+s+t+r)$. As $wt(v)=wt(w)$, it is clear that $s=t$. If $s=t\geq 3$, then we can construct an involution which fixes all basis elements as in the Case 2 of the proof of Proposition 3.13. There are linear codes such that $l_{i}\geq 3$ for some $i\in\{0,1,2\}$, if this is the case we can construct an involution which fixes all basis elements as in the Case 1 of the proof of Proposition 3.13.
It should be mentioned that there are some linear codes with the property that $l_{i}\leq 2$ for all $i$ and $s=t\leq 2$. So we should consider the case $r\geq 1$, that is the number of common zero blocks of $v$ and $w$ is at least one. Let $v|_{\Omega_{i}}=w|_{\Omega_{i}}=000$. The corresponding block of $z$ is $z|_{\Omega_{i}}\in\{000,111\}$. In fact we have a submatrix of $G$ in the form
$\begin{bmatrix}aaa\\ 000\\ 000\\ 000\\ 000\end{bmatrix}$
where $a\in\{0,1\}$. So any transposition on this block (any transposition changing the columns of the above submatrix) fixes all basis elements.
In all cases, we construct an involution in $\mathrm{PAut}(C)$, so such a $C$ does not exist. ∎
Proof of Theorem 1.2.
Follows from Corollary 2.6, Corollary 3.11, and the fact that $\mathrm{PAut}(C)=\mathrm{PAut}(C^{\perp})$ where $C^{\perp}$ is the dual of $C$.
Proof of Theorem 1.3.
Follows from Proposition 3.13 and Proposition 3.14.
Remark 3.15.
Whenever $E_{\sigma}(C)$ is a non distinct $\sigma$-weight code with length $n<30$, we can not construct transpositions as in the proof of Proposition 3.14. For example, for the length $n=6$, let $\sigma=(123)(456)$ and consider a binary linear code with the following generator matrix
$G=\begin{bmatrix}z\\ v\\ w\\ v^{\sigma}\\ w^{\sigma}\end{bmatrix}=\begin{bmatrix}111&000&\\ 110&000&\\ 000&101&\\ 011&000&\\ 000&110&\end{bmatrix}.$
In this case there is no transposition or involution fixing all basis elements. However, the transposition $\alpha=(13)$ satisfies $z^{\alpha}=z$, $v^{\alpha}=v^{\sigma}$, $w^{\alpha}=w$, ${w^{\sigma}}^{\alpha}=w^{\sigma}$. Hence $\alpha\in\mathrm{PAut}(C)$. Moreover calculations in MAGMA give that $\mathrm{PAut}(C)\cong S_{3}\times S_{3}$
Now let $n=9$. Let $\sigma=(123)(456)(789)$ and consider the code generated by the following matrix
$G=\begin{bmatrix}z\\ v\\ w\\ v^{\sigma}\\ w^{\sigma}\end{bmatrix}=\begin{bmatrix}111&000&111&\\ 110&000&101&\\ 000&101&110&\\ 011&000&110&\\ 000&110&011&\end{bmatrix}.$
In this case $\alpha=(12)(56)(79)$ acts on the basis elements as follows
$z^{\alpha}=z,\ v^{\alpha}=v,\ w^{\alpha}=w^{\sigma},\ {(v^{\sigma})}^{\alpha}=v+v^{\sigma},\ {(w^{\sigma})}^{\alpha}=w.$
We observe from the calculations carried out in MAGMA that $\tau=(13)(45)(89)\in\mathrm{PAut}(C)$ and $K=\langle(19),(27),(38)\rangle$ is an elementary abelian $2$-subgroup of order $8$ of ${\rm PAut}(C)$. Let $L=K\rtimes{\langle\sigma\rangle}$ then ${\rm PAut}(C)=L\rtimes{\langle\tau\rangle}$.
Remark 3.16.
The examples in Remark 3.15 show that under the assumption when $n<30$, we can not find a pattern to construct an involution in $\mathrm{PAut}(C)$.
For $5$-dimensional cubic codes of length $n$, we obtain some computational results for the case where $n\in\{6,9,12,15,18,21\}$. These results are listed in Table 1. Indeed, using the construction methods in [7](see Theorem 5.1 and section VI, the case $m=3$), we construct all cubic codes of length $n$, where $n\in\{6,9,12,15,18,21\}$ and determine the order of the permutation automorphism groups of these codes. Moreover, for the case $n=21$, we construct all cubic codes and recognize that the smallest order for the permutation automorphism groups is $12$. For this length we give a lower bound for the number of different cubic codes up to equivalence.
For $n\in\{24,27\}$, with respect to our computer facilities, the calculations seem to take years. Up to date, we get some partial results as follows:
  • For the length $n=24$, with respect to construction methods in [7], there are $8206520925$ cubic codes and up to date, we get $5366$ different cubic codes up to equivalence. Among these cubic codes, the smallest order for the permutation automorphism group is $48$ and the group structure for these permutation automorphism groups of order $48$ is $C_{2}\times S_{4}$.
  • For the length $n=27$, with respect to construction methods in [7], there are $263945834061$ cubic codes and up to date, we get $11440$ different cubic codes up to equivalence. Among these cubic codes, the smallest order for the permutation automorphism is $48$ and the group structure for these permutation automorphism groups of order $48$ is $C_{2}\times S_{4}$.
  • For the length $n\in\{24,27\}$, these partial results support the fact that a cubic code of dimension $5$ and length $n\in\{24,27\}$ has permutation automorphism group of order greater than $3$. All computations are done in SAGEMATH.
Length of Cubic Codes of dimension five 6 9 12 15 18 21
The number of different Cubic Codes up to equivalence 2 15 67 244 765 $\geq 2142$
Minimum order of the permutation automorphism groups 36 12 6 6 6 12
Structure of automorphism groups with minimum order $S_{3}\times S_{3}$ $D_{12}$ $S_{3}$ $S_{3}$ $S_{3}$ $D_{12}$
Table 1.
Acknowledgements
İ. TUVAY was partially supported by the Scientific and Technological Research Council of Turkey (Tübitak) through the research program Bideb-2219.
The numerical calculations reported in this paper were partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).